# Trust-Preserved Human-Robot Shared Autonomy enabled by Bayesian Relational Event Modeling

### v2

@@ I Introduction
As robots continue to find applications in a wide range of real-world scenarios, the next wave of technological progress in the field of robotics appears to be centered around their ability to interact effectively with complex environments and autonomously execute tasks with minimal human supervision [1, 2, 3]. Although maintaining human global control over autonomous systems contributes to increasing the initiative and awareness of dynamic, complex, and interactive environments, the ability of robots to make unaided decisions also grows in significance [4, 5, 6].
On the one hand, the efficiency of the human-robot teams can be compromised due to frequent human intervention [7, 8, 9]. From the human perspective, overseeing and instructing robots, particularly in multi-robot teams where a large group of robots require simultaneous management, can demand a significant commitment of effort. For robots, continuous interruptions from human can cause disturbances that make their operation destabilized, dependent, and slow [10], especially in the presence of communication delays.
On the other hand, the fundamental expectation behind human-robot teaming is that the abilities of robots and humans complement each other, resulting in enhanced performance compared to their individual efforts [11, 12, 13]. Robots can often offer an information advantage by accessing data that are unavailable to human team members [10]. For example, consider a search and rescue (SAR) scenario as illustrated in Fig. 1. The human responder might initially focus on rescuing injured individuals. At the same time, the robot identifies a gas leakage that requires immediate attention to prevent a potential explosion endangering everyone within the scene. In such a situation, the robot is challenged to weigh the importance of two conflicting goals: adhering to the human’s instructions to continue rescuing the victim versus prioritizing putting down the gas leakage. Hence in the best interests of the team as a whole, the robots are expected to make judicious and rational assessments to determine when to adhere to or disregard human directives when conflicts happen.
@@
Shared Autonomy (SA). When teaming up with robots, human’s expectations of robots now have gone beyond merely an “extended hand” to alleviate physical tasks, but an “extended brain” that can support, inform, and even rectify us when necessary [14, 15, 16]. The degree to which the robots autonomously choose to follow human instructions in the light of other objectives can be characterized by their autonomy level [7]. Shared autonomy provides an opportunity to establish a flexible autonomy framework that grants robots the capacity to function across varying autonomy levels [3]. There have been various autonomy adaptation strategies proposed utilizing information extracted from both the human (e.g. intentions, muscle activity, skills) and the environment (e.g. task, constraints) [17, 18, 19, 20, 21, 22].
However, most of those approaches ignore the fact that the autonomous decision-making capabilities of robotic systems can sometimes be unsettling for humans due to various factors, including perceived risks and a lack of trust [23]. Therefore, to achieve more fluent human-robot teaming, concerns regarding the acceptability of non-human entities making decisions autonomously on behalf of human team members also need to be addressed. From a longitudinal perspective on trust development and calibration in human-robot teams [24], the robot’s autonomy level should be well aligned with the human’s trust level to avoid autonomy surprises. Consequently, this paper paid special attention to the maintenance of human trust during the adaptation of the robot’s autonomy level.
Trust Inference. The main challenge that obstructs successful autonomy-trust coordination is the inherent difficulty of quantifying human trust. Numerous researchers have approached this challenge from diverse viewpoints, for example, physiological measures (e.g. skin conductance, eye-tracking), neuroimaging, behavioral observations, and self-reported trust ratings [25, 26, 27]. Those approaches function effectively in the scenarios of proximate human-robot interaction where humans and robots share the same workspace, allowing for ample human observations.
However, the majority of those methods lose their throne in more general real-world scenarios of human-robot teaming where humans and robots are often separated spatially and temporally. For example, in the SAR scenarios, disaster robots are separated from their remote operators in both space and time. In those cases, the only available interaction information is a flow of commands communicated between humans and robots, where none of these events by itself characterizes the team relationships at any point in time. To tackle this challenge, this paper proposed an innovative approach to dynamically infer the underlying drivers of interaction dynamics (e.g. trust) based merely on time-stamped relational events communicated within human-robot teams (e.g. commands).
@@
Trust Inference from Remote Interactions. We propose a novel approach to dynamically infer human trust from remote human-robot interactions by enhancing the relational event modeling (REM) framework [28, 29, 30, 31, 32, 33] with Bayesian learning techniques.
Trust-preserved Shared Autonomy. We devise a trust-preserved shared autonomy strategy that allows the robots to seamlessly adjust their autonomy level while actively establishing, maintaining, and repairing human trust.
@@
II Problem Formulation
@@
We are inspired by a human-robot collaborative SAR scenario where the human and multiple robots jointly perform the search and evacuation of victims from an incident area [34], as illustrated in Fig. 2.
@@
Human-robot communications: Humans and robots can communicate through commands using template sentences.
The human has the option to either acquire the robots’ status or give instructions to the robots.
@@
Furthermore, the robots can also share their understanding of the environment with the human (e.g., “Instruction received. However, I have the following additional environmental information that may be useful for you to refine the decisions. Please resend your decision.”); and provide an explanation of their decisions to the human (e.g., “My priority has been adjusted due to an observed dangerous condition in Building X.”).
@@
Problem Statement: Grounded on the toy example scenario described above, the objective of this paper is to develop a trust-preserved shared autonomy framework that can:
  • (i)
    Dynamically infer the human’s trust in the robots based on merely time-stamped commands communicated between the human and robots;
  • (ii)
    Adaptively coordinate the robot’s autonomy level with the human’s trust level, which facilitates both the completion of the task and the acceptability of the human.
III Bayesian Relational Event Modelling
III-A Relational Event Modelling
Interactions among individuals can be represented as a sequence of relational event, which is defined as “a discrete event generated by a social actor and directed toward one or more targets” [29]. Formally, a relational event is represented by a tuple $e=(i_{e},j_{e},k_{e},w_{e},t_{e})$ containing the sender $i_{e}$, receiver $j_{e}$, type $k_{e}$, weight $w_{e}$, and time $t_{e}$ of an event. The full sequence of $m$ events formulates the history event set $E=\{e_{1},\cdots,e_{m}\}$. Let $A=\{1,\cdots,n\}$ be the set of all $n$ actors, and $D\subseteq A\times A$ the set of all dyads. Events can fall into a set of $K$ discrete classes. For each class $k$, define the weight function $w_{kt}$ as the accumulated weight of all past interaction events [35].
REM is a blend of social sequence and network analysis, which provides a principled approach to investigate the underlying factors that drive the occurrence, rhythm, and speed of individuals interacting with each other over time. REM models every potential pairing event of individuals in terms of rates, where common events have high rates, and rare events have low rates. Taking from the Cox proportional hazards model [36], the rate for a sender $i$ directing an action of type $k$ to receiver $j$ at time $t$ is modeled as a log-linear function of a set of sufficient statistics $S_{p}(i,j,k,G_{t})$ and their corresponding intensity parameters $\theta_{p}$:
where $\boldsymbol{\theta}=[\theta_{1},\cdots,\theta_{P}]^{\intercal}$, and the parameter $\lambda_{0}(t)$ controls the relationship between the rate and time. The simplest case is to assume that this value is constant, which corresponds to the exponential model.
The likelihood function for the full event sequence and timing can be explicitly computed as the probability of each event, multiplied by the probability of the time between events, conditional on the entire realized history. Leaving out the time-dependent parameter $\lambda_{0}(t)$, the partial likelihood function is given by
$\begin{split}&f(E;G_{t},\boldsymbol{\theta})=\prod_{e\in E}\lambda_{i_{e}j_{e}k_{e}}(t_{e}|G_{t_{e-1}},\boldsymbol{\theta})\times\\ &\exp{\left(-\Delta t_{e}\sum_{k=1,\cdots,K}\sum_{(i,j)\in D}\lambda_{ijk}(t_{e}|G_{t_{e-1}},\boldsymbol{\theta})\right)},\end{split}$ (1)
where $\Delta t_{e}=t_{e}-t_{e-1}$ is the time elapsed between events. This expression represents the likelihood of every event in the sequence, along with the conditional likelihood that no other events occurred in the period between observed events.
If the relational event data is simply ordinal, i.e. there is no specific timing data available, the likelihood function can be reduced to a product of multinomial probabilities derived from each event rate
$f(E;G_{t},\boldsymbol{\theta})=\prod_{e\in E}\frac{\lambda_{i_{e}j_{e}k_{e}}(t_{e}|G_{t_{e-1}},\boldsymbol{\theta})}{\sum_{k=1,\cdots,K}\sum_{(i,j)\in D}\lambda_{ijk}(t_{e}|G_{t_{e-1}},\boldsymbol{\theta})}.$
III-B Sufficient Statistics
The greatest utility of REM lies in the wide range of sufficient statistics that can be derived from event sequences. These measures are numerical representations of specific interaction patterns, similar to those encoded in exponential random graph models (ERGMs) or stochastic actor-oriented models [37]. That is, like how conceptual models are translated into statistical models, sufficient statistics are the operationalizations of model parameters including both endogenous and exogenous factors.
Exogenous factors refer to characteristics outside the relational event history, including individual attributes, relational attributes, as well as environmental contextual information beyond the scope of the interaction system. For example, sometimes the likelihood of a group interaction occurring is due to some individual attributes of either the sender or receiver. An extroverted individual might send messages at a higher rate than somebody who is more introverted. Relational attributes refer to different kinds of ties individuals might have with one another like affinity (e.g., friendship, trust), flow (e.g., other forms of sending messages), representational (e.g., endorsements), and semantic (e.g., shared interpretations) ties [38]. Environmental information may encompass more readily measurable entities like the nature of the task, restrictions on communication channels, or availability of resources.
@@ III-B Sufficient Statistics
Fig. 3 describes some sufficient statistics used to model the group communications for the collaborative SAR scenario investigated in this paper. For endogenous factors, we include inertia and reciprocity. For exogenous factors, we include the human’s individual attribute of command and the relational attribute of trust between the human and the robots. Those sufficient statistics are by no means exhaustive, and various other structures may be encoded to capture specific behavioral patterns if needed.
@@
III-C Bayesian Relational Event Modelling
To dynamically infer the evolution of human trust, we propose to advance REM into an adaptive framework taking advantage of Bayesian learning. This framework is composed of two stages: 1) offline learning for team properties grounding, and 2) online learning for dynamic attributes inference.
Team properties grounding. The sign and magnitude of each element in the coefficient parameter $\boldsymbol{\theta}$ determines how influential a particular network effect is regarding the generation of relational events across team members. Therefore, the first stage aims to estimate the coefficient parameter $\boldsymbol{\theta}$ to ground the specific properties of the team to be investigated. In this stage, a period of interaction history $E_{0}$ within the team is collected, as well as the initial attribute information $G_{0}$ that can be obtained by a post-process survey. Then the coefficient estimates for relational event models – ordinal or temporal – are derived from the solution of the maximum likelihood problem:
@@ III-C Bayesian Relational Event Modelling
Common optimization techniques such as Newton-Rhapson are well-suited to solving this maximization problem.
Dynamic attributes inference. Based on the grounded properties $\boldsymbol{\theta}^{\ast}$ for a specific team, the next stage is to deploy a receding horizon approach to dynamically infer the time-varying attributes $G_{t}$ in the team (e.g., trust), as illustrated in Fig. 4. We assume that within each observation interval $T_{k}$ between two discrete observation time instant $t_{k-1}$ and $t_{k}$, the attributes $G_{t}$ remain constant. And the observed relational events history within the observation interval $T_{k}$ is recorded as $E_{k}$. Define $\pi_{k}(G_{t})=P(G_{t}|E_{1:k},\boldsymbol{\theta}^{\ast})$ as the posterior distribution of $G_{t}$ after $k$-th observation intervals. According to the Bayes’ Theorem,
@@
where the likelihood function $f(E_{k};G_{t},\boldsymbol{\theta}^{\ast})$ can be calculated according to (1), and the prior distribution $\pi_{0}(G_{t})$ offers the opportunity to take the initial attribute information $G_{0}$ collected in the previous stage into consideration.
IV Trust-Preserved Shared Autonomy
IV-A Quantification of Trust and Autonomy Levels
Following [39], we define trust as “the willingness of a party to be vulnerable to the actions of another party based on the expectation that the other will perform a particular action important to the trustor, irrespective of the ability to monitor or control that other party”. By including the human trust $\beta$ into the time-varying attributes $G_{t}$, we can dynamically infer the evolution of trust based on Bayesian REM. Superior to some point estimation approaches, e.g., maximum likelihood estimation (MLE), the utilization of Bayesian learning enables us to maintain the complete statistic description of trust through a probability distribution $\pi(\beta)$, which offers greater flexibility in quantifying trust level tailored to specific scenarios. For example, the trust level $L_{\beta}$ in general can be determined by taking expectation of the estimated trust distribution: $L_{\beta}=\mathbb{E}_{\pi(\beta)}\beta$. However, in some safety-critical scenarios, we may prefer to take a risk-averse perspective with certain risk measures: $L_{\beta}=\mathcal{R}_{\pi(\beta)}\beta$, such as mean-variance, Value at Risk (VaR), and Conditional Value at Risk (CVaR) (cf. [40]).
Level of robot autonomy, ranging from teleoperation to fully autonomous systems, defines the extent to which a system can carry out its own processes and operations without external control. There have been different criteria of assessing the autonomy level, for example, the allocation between functions, or the amount of time that a person can neglect the robot [7]. In the context of this paper, we quantify the autonomy level $L_{\alpha}$ of the robot as $L_{\alpha}=1-\alpha$, where $\alpha$ is the probability that the robot follows the human’s instruction when contradicting its own reasoning.
IV-B Coordination between Trust and Autonomy Levels
After formally quantifying the human trust level and robot autonomy level, the immediate question that arises is: How to effectively adjust the robot autonomy level according to the evolving human trust level? In general, the more trust humans place on robots, the higher level of autonomy they tend to accept. Therefore, an intuitive adaptive strategy is to passively choose the autonomy level that best matches the estimated trust level: $L_{\alpha}\propto L_{\beta}$, as illustrated in Fig. 5.
However, this reactive adjustment may impede the effectiveness and efficiency of the human-robot team and fail to take full advantage of the robot’s capabilities. For example, when humans maintain a lower level of trust in the robots, disuse or micromanagement may happen, which can result in suboptimal solutions to both task completion and increased workload for individual members.

### v3

@@ I Introduction
As robots continue to find applications in a wide range of real-world scenarios, the next wave of technological progress in the field of robotics appears to be centered around their ability to interact effectively with complex environments and autonomously execute tasks with minimal human supervision [Selvaggio2021AutonomySurvey]. Although maintaining human global control over autonomous systems contributes to increasing the initiative and awareness of dynamic, complex, and interactive environments, the ability of robots to make unaided decisions also grows in significance [Liu2022CoordinatingProficiencies].
On the one hand, the efficiency of the human-robot teams can be compromised due to frequent human intervention [Beer2014TowardInteraction, Franchi2012BilateralTopology]. From the human perspective, overseeing and instructing robots, particularly in multi-robot teams where a large group of robots require simultaneous management, can demand a significant commitment of effort. For robots, continuous interruptions from human can cause disturbances that make their operation destabilized, dependent, and slow [Schermerhorn2009DynamicTask], especially in the presence of communication delays.
On the other hand, the fundamental expectation behind human-robot teaming is that the abilities of robots and humans complement each other, resulting in enhanced performance compared to their individual efforts [Angleraud2021CoordinatingCommands]. For example, robots can often offer an information advantage by accessing data that are unavailable to human team members [Schermerhorn2009DynamicTask]. Consider a human-robot collaborative search and rescue (SAR) scenario as illustrated in Fig. 1. The human responder might initially focus on rescuing injured individuals. At the same time, the robot identifies a gas leakage that requires immediate attention to prevent a potential explosion endangering everyone within the scene. In such a situation, the robot is challenged to weigh the importance of two conflicting goals: adhering to the human’s instructions to continue rescuing the victim versus prioritizing putting down the gas leakage. Hence in the best interests of the team as a whole, the robots are expected to make judicious and rational assessments to determine when to adhere to or disregard human directives when conflicts happen.
@@
The degree to which the robots autonomously choose to follow human instructions in the light of other objectives can be characterized by their autonomy level [Beer2014TowardInteraction]. Shared autonomy [Mostafa2019AdjustableReview], also referred to as adaptive autonomy, provides an opportunity to establish a flexible autonomy framework that grants robots the capacity to function across varying autonomy levels. However, the autonomous decision-making capabilities of robotic systems can sometimes be unsettling for humans due to various factors, including perceived risks and a lack of trust [Desai2009CreatingSystems]. Therefore, achieving fluent human-robot teaming mandates addressing concerns regarding the acceptability of non-human entities making decisions autonomously on behalf of human team members. As such, this paper pays special attention to maintaining human trust during the adaptation of the robot’s autonomy level.
Motivated by the shared properties and constraints of real-world human-robot teaming where humans and robots may be spatially and temporally separated, we investigate a simple but representative scenario of human-robot collaborative SAR [Schoonderwoerd2022DesignTask]. In this scenario, humans and robots have complementary capabilities, and tasks are dependent on sequences, requiring collaboration among different team members. In addition, since the robots are separated from their remote human operators in both space and time, their communication are restricted to merely time-stamped commands communicated between them.
To captivate trust-preserved shared autonomy within such challenging circumstances, the biggest obstacle comes from the inherent difficulty of inferring human trust from solely a flow of commands communicated within the human-robot team, where none of these events by themselves characterize the team relationships at any point in time. To tackle this challenge, we propose a novel approach to dynamically infer the underlying drivers of interaction dynamics (e.g. trust) from merely time-stamped relational events (e.g. commands) by enhancing the relational event modeling (REM) framework [Schecter2021TheModel, Mulder2019ModelingAnalysis] with Bayesian learning techniques. The human trust inferred online is thus an evolving process, which allows us to take a longitudinal perspective on trust development and calibration in human-robot teams [DeVisser2020TowardsTeams]. Leveraging the punctuated equilibrium model (PEM) [Sabherwal2001TheModel], we developed a trust-preserved shared autonomy strategy that allows the robots to seamlessly adjust their autonomy levels while actively establishing, maintaining, and repairing human trust.
@@
Trust Inference from Remote Interactions. We propose an innovative approach to dynamically infer human trust based on remote human-robot interactions.
Trust-preserved Shared Autonomy. We devise a trust-preserved shared autonomy strategy that maintains human trust during the adaptation of robot’s autonomy level.
@@
II Related Work
Trust Inference. The priority of trustworthy human-robot teaming is to capture a relatively accurate understanding of human trust. In the field of human-robot interaction, numerous researchers have approached this challenge from diverse viewpoints, for example, physiological measures (e.g. skin conductance, eye-tracking), neuroimaging, behavioral observations, and self-reported trust ratings [Khalid2020DeterminantsPredicting, Guo2020ModelingApproach, Zheng2018Human-RobotSystems]. Those approaches function effectively in the scenarios of proximate human-robot interaction where humans and robots share the same workspace, allowing for ample human observations. However, the majority of those methods lose their throne in more general real-world scenarios of human-robot teaming where humans and robots are separated spatially and temporally. This paper opens new possibilities of rigorously inferring human trust from remote human-robot interactions, which possesses the potential to be applied to more general and complicated real-world human-robot teaming scenarios.
Shared Autonomy (SA). Shared autonomy approaches allow the robot to seamlessly adapt its autonomy level based on its own understanding of the human actions/intentions and of the surrounding environment. There have been various autonomy adaptation strategies proposed utilizing information extracted from both the human (e.g. intentions, muscle activity, skills) and the environment (e.g. task, constraints) [Li2015ContinuousControl, Milliken2017ModelingAutonomy, Enayati2018Skill-basedTracking, Jain2020ProbabilisticRobotics]. Although effective to a certain extent, few of those approaches take into account an essential fact that real-world human-robot teaming is a long-term and evolutionary process. Instead, this paper takes a longitudinal perspective on trust development and calibration in human-robot teams, and devises a trust-preserved shared autonomy strategy that actively maintains human trust during the adaptation of robot’s autonomy level.
III Background on Relational Event Modelling
III-A Relational Event Modelling
Interactions among individuals can be represented as a sequence of relational event, which is defined as “a discrete event generated by a social actor and directed toward one or more targets” [Butts2017ADynamics]. Formally, a relational event is represented by a tuple $e=(i_{e},j_{e},k_{e},w_{e},t_{e})$ containing the sender $i_{e}$, receiver $j_{e}$, type $k_{e}$, weight $w_{e}$, and time $t_{e}$ of an event. Let $A=\{1,\cdots,n\}$ be the set of all $n$ actors, and $D\subseteq A\times A$ the set of all dyads. Events can fall into a set of $K$ discrete classes. For each class $k$, define the weight function $w_{kt}$ as the accumulated weight of all past interaction events [Brandes2009NetworksData].
The full sequence of $m$ events formulates the history event set $E=\{e_{1},\cdots,e_{m}\}$. The occurrence of future relational events can be influenced by both the history of past events $E$ and the attributes within the network $G_{t}$. REM is a statistical technique that blends social sequence and network analysis, which provides a principled approach to investigate the underlying factors that drive the occurrence, rhythm, and speed of individuals interacting with each other over time. REM models every potential pairing event of individuals in terms of rates, where common events have high rates, and rare events have low rates.
Following the event history approach [Blossfeld2001TechniquesAnalysis], the probability of the sequence and timing of events can be defined by the hazard rate and survival function for each possible action. The hazard rate can be interpreted as the instantaneous likelihood of an event occurring, given that it has not yet occurred. The survival function can be interpreted as the likelihood of an event occurring. Survival models relate the time that passes, before some event occurs, to one or more covariates that may be associated with that quantity of time. Proportional hazards models are a class of survival models assuming that, the unique effect of a unit increase in a covariate is multiplicative with respect to the hazard rate, which allows us to estimate the effect parameters without any consideration of the full hazard function. Therefore, REM utilizes the Cox proportional hazards model [Cox1972RegressionLifeTables], where the hazard rate for a sender $i$ directing an action of type $k$ to receiver $j$ at time $t$ is modeled as a log-linear function of a set of sufficient statistics $S_{p}(i,j,k,G_{t})$ and their corresponding intensity parameters $\theta_{p}$:
where $\boldsymbol{\theta}=[\theta_{1},\cdots,\theta_{P}]^{\intercal}$, and $p\in\{1,\cdots,P\}$ is the index of the sufficient statistics, and the parameter $\lambda_{0}(t)$ describes how the risk of event per time unit changes over time at baseline levels of covariates.
The likelihood function for the full event sequence and timing can be explicitly computed by combining the hazard rate for each realized event with the survival function for all possible events. When assuming time-invariant baseline conditions, $\lambda_{0}(t)$ can be chosen as a constant, which gives the partial likelihood function as
$\begin{split}&f(E;G_{t},\boldsymbol{\theta})=C(\boldsymbol{\theta})\prod_{e\in E}\lambda_{i_{e}j_{e}k_{e}}(t_{e}|G_{t_{e-1}},\boldsymbol{\theta})\times\\ &\exp{\left(-\Delta t_{e}\sum_{k=1,\cdots,K}\sum_{(i,j)\in D}\lambda_{ijk}(t_{e}|G_{t_{e-1}},\boldsymbol{\theta})\right)},\end{split}$ (1)
where $\Delta t_{e}=t_{e}-t_{e-1}$ is the time elapsed between the current event $e$ and the previous event e1e-1, and $C(\boldsymbol{\theta})$ is a normalization factor that makes this partial likelihood function well-defined. This expression represents the likelihood of every event in the sequence, along with the conditional likelihood that no other events occurred in the period between observed events.
III-B Sufficient Statistics
The greatest utility of REM lies in the wide range of sufficient statistics that can be derived from event sequences. These measures are numerical representations of specific interaction patterns, similar to those encoded in exponential random graph models (ERGMs) or stochastic actor-oriented models [Lusher2013ExponentialNetworks]. That is, like how conceptual models are translated into statistical models, sufficient statistics are the operationalizations of model parameters including both endogenous and exogenous factors.
Exogenous factors refer to characteristics outside the relational event history, including individual attributes, relational attributes, as well as environmental contextual information beyond the scope of the interaction system. For example, sometimes the likelihood of a group interaction occurring is due to some individual attributes of either the sender or receiver. An extroverted individual might send messages at a higher rate than somebody who is more introverted. Relational attributes refer to different kinds of ties individuals might have with one another like affinity (e.g., friendship, trust), flow (e.g., other forms of sending messages), representational (e.g., endorsements), and semantic (e.g., shared interpretations) ties [Shumate2013ANetworks]. Environmental information may encompass more readily measurable entities like the nature of the task, restrictions on communication channels, or availability of resources.
IV Problem Formulation
@@
We are inspired by a human-robot collaborative SAR scenario where the human and multiple robots jointly perform the search and evacuation of victims from an incident area [Schoonderwoerd2022DesignTask], as illustrated in Fig. 2.
@@
Human-robot communications: Humans and robots can communicate through commands using template sentences. The human has the option to either acquire the robots’ status or give instructions to the robots.
@@
Problem Statement: Based on the toy example scenario described above, the objective of this paper is to develop a trust-preserved shared autonomy framework that can: 1) dynamically infer the human’s trust in the robots based on merely time-stamped commands communicated between them; and 2) adaptively coordinate the robot’s autonomy level with the human’s trust level, which facilitates both the completion of the task and the acceptability of the human.
V Dynamic Trust Inference
V-A Quantification of Trust
Following [Mayer1995AnTrust], we define trust as “the willingness of a party to be vulnerable to the actions of another party based on the expectation that the other will perform a particular action important to the trustor, irrespective of the ability to monitor or control that other party”. It is an essential component that is in general included in the time-varying attributes $G_{t}$. Since we are paying special attention to the trust between team members, we assume that the other attributes are known and static. For simplicity, in the rest of this paper, we denote $G_{t}$ to represent the unknown and dynamic trust that human places on the robots.
V-B Sufficient Statistics
@@
In the context of collaborative SAR scenario investigated in this paper, we utilize the sufficient statistics described in Fig. 3 to model the group communications. For endogenous factors, we include inertia and reciprocity. For exogenous factors, we include the human’s individual attribute of command and the relational attribute of trust between the human and the robots. Those sufficient statistics are by no means exhaustive, and various other structures may be encoded to capture specific behavioral patterns if needed.
@@
V-C Bayesian Relational Event Modelling
To dynamically infer the evolution of human trust, we propose to advance REM into an adaptive framework taking advantage of Bayesian learning. This framework is composed of two stages: 1) offline learning for team properties grounding, and 2) online learning for dynamic trust inference.
(1) Team properties grounding. The sign and magnitude of each element in the coefficient parameter $\boldsymbol{\theta}$ determines how influential a particular network effect is regarding the generation of relational events across team members. Therefore, the first stage aims to estimate the coefficient parameter $\boldsymbol{\theta}$ to ground the specific properties of the team to be investigated. In this stage, a period of interaction history $E_{0}$ within the team is collected, as well as the initial trust information $G_{0}$ that can be obtained by a post-process survey. Then the coefficient estimates for relational event models are derived from the solution of the maximum likelihood problem:
@@ V-C Bayesian Relational Event Modelling
Common optimization techniques such as Newton-Raphson are well-suited to solving this maximization problem.
(2) Dynamic trust inference. Based on the grounded properties $\boldsymbol{\theta}^{\ast}$ for a specific team, the next stage is to deploy a receding horizon approach to dynamically infer the time-varying trust that humans place on the robots, as illustrated in Fig. 4. We assume that within each observation interval $T_{k}$ between two discrete observation time instant $t_{k-1}$ and $t_{k}$, the trust $G_{t}$ remains constant. And the observed relational events history within the observation interval $T_{k}$ is recorded as $E_{k}$. Define $\pi_{k}(G_{t})=P(G_{t}|E_{1:k},\boldsymbol{\theta}^{\ast})$ as the posterior distribution of $G_{t}$ after $k$-th observation intervals. According to the Bayes’ Theorem,
@@
where the likelihood function $f(E_{k};G_{t},\boldsymbol{\theta}^{\ast})$ can be calculated according to (1), and the prior distribution $\pi_{0}(G_{t})$ offers the opportunity to take the initial trust information $G_{0}$ collected in the previous stage into consideration.
Superior to some point estimation approaches, e.g., maximum likelihood estimation (MLE), the utilization of Bayesian learning enables us to maintain the complete statistic description of trust through a probability distribution $\pi(G_{t})$, which offers greater flexibility in quantifying trust level tailored to specific scenarios. For example, the trust level $L_{\text{trust}}$ in general can be determined by taking expectation of the estimated trust distribution: $L_{\text{trust}}=\mathbb{E}_{\pi(G_{t})}G_{t}$. However, in some safety-critical scenarios, we may prefer to take a risk-averse perspective with certain risk measures: $L_{\text{trust}}=\mathcal{R}_{\pi(G_{t})}G_{t}$, such as mean-variance, Value at Risk (VaR), and Conditional Value at Risk (CVaR) (cf. [Wu2018AAsymptotics]).
VI Trust-Preserved Shared Autonomy
VI-A Quantification of Autonomy
Level of robot autonomy, ranging from teleoperation to fully autonomous systems, defines the extent to which a system can carry out its own processes and operations without external control. There have been different criteria for assessing the autonomy level, for example, the allocation between functions, or the amount of time that a person can neglect the robot [Beer2014TowardInteraction]. In the context of this paper, we quantify the autonomy level $L_{\text{autonomy}}$ of the robot as $L_{\text{autonomy}}=1-\alpha$, where $\alpha$ is the probability that the robot follows the human’s instruction when contradicting its own reasoning.
VI-B Coordination between Trust and Autonomy
Based on the inferred dynamics of trust level, the immediate question that arises is: How to effectively adjust the robot autonomy level according to the evolving human trust level? In general, the more trust humans place on robots, the higher level of autonomy they tend to accept. Therefore, an intuitive adaptive strategy is to passively choose the autonomy level that best matches the estimated trust level: $L_{\text{autonomy}}\propto L_{\text{trust}}$, as illustrated in Fig. 5. However, this reactive adjustment may impede the effectiveness and efficiency of the human-robot team and fail to take full advantage of the robot’s capabilities. For example, when humans maintain a lower level of trust in the robots, disuse or micromanagement may happen, which may impede the completion of task and increase the workload for human operators.